# Artificial Intelligence in Elementary STEM Education: A Systematic Review of Current Applications and Future Challenges

### v1

@@ 3.14.5 Privacy-Preserving Algorithmic Techniques and Data Governance
Sensor and feature selection priorities, as detailed above, emphasize non-identifying modalities (low-resolution thermal, capacitive, aggregated logs, environmental sensors) while limiting adoption of high-identifiability sensors (RGB cameras, continuous audio) unless strict safeguards and pedagogical necessity justify their use [48, 62, 63].
@@
On-device and edge processing with data minimization perform initial signal processing, feature extraction, and preliminary inference locally within the classroom computing environment, transmitting only aggregated, de-identified, or sufficiently obfuscated feature representations needed for downstream teacher dashboards and learning analytics [56, 68, 69]. For example, thermal array data undergoes CNN-based occupancy classification on the classroom edge node, outputting zone occupancy counts and coarse activity labels (sitting, standing, walking) while raw thermal frames are discarded after processing. This architectural pattern reduces raw data exfiltration risk, minimizes storage requirements, and enables real-time teacher feedback with millisecond-range latency [56, 64].
@@
Feature-level obfuscation and camouflage learning techniques apply controlled perturbation or masking to sensitive feature values before transmission or storage, preserving sufficient statistical utility for aggregate analysis while reducing the risk of reconstructing identifying information from shared features [68]. Methods proposed for resource-constrained IoT devices in AmI contexts include additive noise injection calibrated to preserve classification accuracy, feature subsampling that omits fine-grained temporal detail, and differential privacy mechanisms providing formal privacy guarantees with quantified privacy budgets.
@@
Provenance and metadata management maintains annotated data provenance tracking the origin, processing history, consent status, and retention policies associated with each data record or feature set [70]. Provenance-aware metadata frameworks enable auditing of data flows, support conditional consent enforcement (e.g., data authorized only for teacher dashboard display but not for external research sharing), and facilitate regulatory compliance by documenting data lineage. Semantic metadata and ontology-based annotation schemes developed for smart buildings and cyber-physical systems provide templates adaptable to educational contexts [71, 72, 73].
@@
Security mechanisms including encrypted communication channels, authenticated key management, intrusion detection systems (IDS) tailored to IoT/AmI device profiles, and secure firmware provisioning protect against unauthorized data access, sensor spoofing, and network eavesdropping [74, 75, 76]. Lightweight cryptographic protocols suitable for resource-constrained sensors, blockchain-based audit logs for immutable provenance trails, and anomaly detection monitoring for unusual data access patterns constitute a defense-in-depth security architecture necessary for protecting sensitive classroom data.
@@
Policy frameworks and standards alignment leverage guidelines and interoperability practices proven in Active Assisted Living (AAL) and health AmI contexts, including policy templates for consent, retention, data sharing agreements, and cross-institutional governance structures [77, 71, 72]. Formal policy mechanisms for context-sensitive data acquisition and access gating developed for distributed cyber-physical systems can be repurposed for school district networks, specifying role-based access controls, purpose limitations, and automated policy enforcement [78].
@@
Together, these techniques form a comprehensive privacy architecture layering hardware selection, local processing, algorithmic obfuscation, cryptographic protection, provenance tracking, and institutional governance to minimize identifiability while preserving pedagogically actionable signals [68, 70, 56, 74, 76].
@@ 3.14.6 Ethical Frameworks, Stakeholder Engagement, and Participatory Design
Early and iterative stakeholder engagement involves end users (students with developmentally appropriate assent processes), caregivers (parents/guardians providing informed consent), teachers (implementing technology and interpreting analytics), administrators (ensuring policy compliance and resource allocation), and independent ethics advisors throughout design cycles to surface values, acceptability thresholds, and perceived risks [49, 53, 54]. MMLA with young children requires parent/guardian informed consent documents that clearly explain sensor modalities, data uses, retention policies, access controls, and opt-out procedures in accessible language, complemented by child-appropriate assent processes using visual aids and age-appropriate explanations respecting children’s developing autonomy [47, 49].
Values-oriented and participatory design methodologies, adapted from AmI healthcare platform development, align technology architectures with pedagogical aims, community norms, and ethical values through structured design activities [53, 79, 49]. Co-design workshops with teachers and parents identify acceptable use cases, unacceptable surveillance practices, and preferred transparency mechanisms, while child-centered design activities (drawing desired classroom technologies, role-playing sensor scenarios) elicit children’s perspectives on privacy and technology in their learning environments.
Risk assessment and duty of care require formal privacy, psychosocial, and pedagogical risk assessments before deployment, recognizing that continuous behavioral monitoring may alter classroom relationships, create surveillance anxiety, produce stigmatizing inferences, or generate data vulnerable to misuse [49, 54, 80]. AmI healthcare literature highlights how ambient sensing can subtly shift power dynamics and trust relationships, underscoring the need for ongoing risk monitoring, incident response protocols, and mechanisms for students/families to raise concerns without penalty.
Legal and institutional alignment ensures technical architectures comply with applicable education privacy laws (FERPA in U.S. contexts, GDPR in European contexts), district data governance policies, and ethical research standards when MMLA systems serve dual educational and research purposes [80, 77]. Clear policies specifying data ownership (student/family, school district, technology vendor), retention limits (automated deletion schedules, prohibition on indefinite retention), permissible uses (restricted to specified pedagogical or research purposes), and prohibited uses (commercial profiling, predictive discipline, third-party sale) provide institutional safeguards complementing technical protections.
@@ 3.14.7 User Experience and Teacher Dashboard Design
UX evaluation and acceptability testing are essential components of AmI deployment, with standardized evaluation instruments available for measuring perceived usefulness, trust, privacy concerns, cognitive load, and unintended consequences [79, 53]. For elementary MMLA interfaces, design priorities emphasize presenting aggregated, actionable indicators to teachers (class-level engagement trends, zone-based attention heatmaps, activity-type participation patterns) rather than granular individual student judgments, unless explicit consent, demonstrated pedagogical necessity, and strict governance safeguards justify finer-grained displays [47, 53, 49].
Teacher dashboards should support pedagogical decision-making without overwhelming educators with excessive detail or creating punitive surveillance atmospheres. Effective designs provide at-a-glance class climate indicators (overall engagement level, attention distribution across learning zones, participation balance), customizable alerts for situations warranting intervention (prolonged disengagement cluster, environmental conditions degrading below thresholds), and historical trend visualizations supporting reflection on instructional effectiveness [47]. Critically, dashboards must avoid reducing complex learning processes to simplistic numeric scores or traffic-light color schemes that encourage superficial interpretations or stigmatizing labels.
@@ 3.14.8 Operational Blueprint: Stepwise Implementation Checklist
Phase 1 - Co-design and needs analysis: Convene teachers, parents, school administrators, and child representatives using structured participatory design methods; define pedagogical objectives, acceptable intervention modes, and privacy boundaries grounded in values-oriented design frameworks [47, 53, 49].
@@
Phase 2 - Risk assessment and policy specification: Conduct formal privacy impact assessment, psychosocial risk analysis, and pedagogical necessity review; specify data retention schedules, consent procedures, access control matrices, incident response protocols, and audit mechanisms modeled on AAL and AmI healthcare guidance [77, 49, 54].
@@
Phase 3 - Sensor selection and pilot hardware: Choose least-identifying sensor modalities meeting pedagogical requirements (e.g., low-resolution thermal arrays for occupancy, capacitive floors for trajectories, environmental sensors for context, aggregated device logs for participation) while avoiding routine RGB camera or continuous audio use unless explicitly justified and safeguarded [48, 55, 58, 62].
@@
Phase 4 - Edge-centric architecture and secure networking: Design classroom edge computing nodes for local feature extraction and obfuscation, implement secure key management and encrypted communication channels, deploy intrusion detection monitoring, and configure network policies to transmit only aggregated features as specified by governance policies [56, 74, 76].
@@
Phase 5 - Algorithmic design and privacy measures: Implement sensor-appropriate machine learning models (CNNs for spatial sensor fusion, LSTMs for temporal trajectory modeling) with feature-level obfuscation techniques, differential privacy mechanisms where applicable, and minimal necessary data retention enforced through automated deletion schedules [48, 66, 68, 61].
@@
Phase 6 - Provenance, auditing, and interoperability: Tag all processed outputs with provenance metadata documenting origin, processing steps, consent status, and retention policies to enable auditing and regulatory compliance; adopt interoperable metadata schemas and ontology practices to support potential future integration while preserving policy-driven access controls [70, 71, 72].
@@
Phase 7 - Pilot evaluation and UX testing: Deploy system in small-scale classroom pilots using standardized UX questionnaires and pedagogical outcome measures; evaluate teacher workflow impacts, psychosocial effects on students and classroom climate, technical reliability, and model performance; iterate design based on stakeholder feedback [79, 53, 47].
@@
Phase 8 - Scale-up governance and continuous monitoring: Establish ongoing risk monitoring processes, model maintenance and retraining schedules to address concept drift, stakeholder feedback channels for reporting concerns, periodic audits of data access logs and policy compliance, and governance committee oversight with teacher, parent, and administrator representation [67, 49, 81].
@@ 3.14.9 Minimal-Risk Reference Configuration
Sensors: Overhead low-resolution thermal array (8x8 or 16x16 pixels) for non-identifying occupancy and zone activity classification [48]; capacitive proximity floor tiles for trajectory extraction without visual capture [55]; ambient environmental sensors (CO2, noise level, illuminance) for contextual class-level activity proxies [58]; aggregated classroom device event logs from tablets/interactive displays for participation metrics [56].
Architecture: Classroom edge computing node executes CNN-based occupancy inference and LSTM-based short-horizon trajectory summarization, applies feature-level obfuscation via camouflage learning techniques, and exports only zone-aggregated engagement scores and participation summaries to teacher dashboard; raw sensor streams retained transiently on edge device (minutes to hours) for debugging purposes per policy, then automatically purged [48, 66, 68, 56].
Governance: Informed parental/guardian consent with child assent; teacher dashboard limited to class/zone-level aggregates without individual student tracking; provenance labels attached to all exported data summaries documenting consent status and permissible uses; regular independent audits of data access logs; family opt-out pathway allowing non-participation without penalty [49, 53, 70].
This configuration follows AmI precedents for non-identifying sensing and edge-centric processing to preserve privacy while enabling pedagogically valuable real-time engagement awareness [48, 55, 56, 68].
@@ 3.14.10 Persistent Challenges and Research Gaps
Limited elementary evidence base: Systematic reviews of MMLA with young children identify comparatively small sample sizes, methodological heterogeneity, and limited longitudinal validation, indicating urgent need for domain-specific validation studies employing child-centered evaluation frameworks and developmentally appropriate measures [47].
Model interpretability and fairness: Adoption of computer vision and affective computing in educational contexts demands heightened caution regarding over-interpretation of algorithmic inferences, fairness across demographic subgroups, and risks of misclassification producing stigmatizing labels or inappropriate interventions; medical CV literature emphasizes the necessity of interpretability, domain-specific validation, and human oversight when algorithms influence consequential decisions about children [50].
System sustainability and model lifecycle: HAR and AmI models exhibit finite operational lifespans due to concept drift, environmental changes, and evolving classroom practices, requiring robust maintenance plans, continuous performance monitoring, and periodic retraining protocols to sustain accuracy and prevent silent degradation [67].
Interoperability and metadata standards: While ontologies and metadata schemas exist for smart buildings and cyber-physical systems, educational technology ecosystems lack widely adopted standards for MMLA data exchange, creating barriers to cross-institutional research, longitudinal student tracking across grade transitions, and integration with existing learning management systems [71, 72, 73].
IoT security and attack surfaces: Classroom IoT sensors present unique security challenges including physical tampering risks, wireless network vulnerabilities, and resource constraints limiting cryptographic capabilities, demanding standardized secure provisioning frameworks and tailored intrusion detection systems [76, 74, 75].

### v2

@@ 3.14.5 Privacy-Preserving Algorithmic Techniques and Data Governance
Sensor and feature selection priorities, as detailed above, emphasize non-identifying modalities (low-resolution thermal, capacitive, aggregated logs, environmental sensors) while limiting adoption of high-identifiability sensors (RGB cameras, continuous audio) unless strict safeguards and pedagogical necessity justify their use [DOI_3390_s21041036, DOI_3390_electronics10050559, DOI_25046_aj050439].
@@
On-device and edge processing with data minimization perform initial signal processing, feature extraction, and preliminary inference locally within the classroom computing environment, transmitting only aggregated, de-identified, or sufficiently obfuscated feature representations needed for downstream teacher dashboards and learning analytics [DOI_33166_aetic_2020_05_001, DOI_1109_tmc_2021_3092271, DOI_3390_cryptography8010005]. For example, thermal array data undergoes CNN-based occupancy classification on the classroom edge node, outputting zone occupancy counts and coarse activity labels (sitting, standing, walking) while raw thermal frames are discarded after processing. This architectural pattern reduces raw data exfiltration risk, minimizes storage requirements, and enables real-time teacher feedback with millisecond-range latency [DOI_33166_aetic_2020_05_001, DOI_1097_qmh_0000000000000423].
@@
Feature-level obfuscation and camouflage learning techniques apply controlled perturbation or masking to sensitive feature values before transmission or storage, preserving sufficient statistical utility for aggregate analysis while reducing the risk of reconstructing identifying information from shared features [DOI_1109_tmc_2021_3092271]. Methods proposed for resource-constrained IoT devices in AmI contexts include additive noise injection calibrated to preserve classification accuracy, feature subsampling that omits fine-grained temporal detail, and differential privacy mechanisms providing formal privacy guarantees with quantified privacy budgets.
@@
Provenance and metadata management maintains annotated data provenance tracking the origin, processing history, consent status, and retention policies associated with each data record or feature set [DOI_1007_s41019-020-00118-0]. Provenance-aware metadata frameworks enable auditing of data flows, support conditional consent enforcement (e.g., data authorized only for teacher dashboard display but not for external research sharing), and facilitate regulatory compliance by documenting data lineage. Semantic metadata and ontology-based annotation schemes developed for smart buildings and cyber-physical systems provide templates adaptable to educational contexts [DOI_3390_smartcities5030053, DOI_3390_en14072024, DOI_3390_su13105578].
@@
Security mechanisms including encrypted communication channels, authenticated key management, intrusion detection systems (IDS) tailored to IoT/AmI device profiles, and secure firmware provisioning protect against unauthorized data access, sensor spoofing, and network eavesdropping [DOI_1109_access_2020_2995917, DOI_1109_access_2022_3179418, DOI_1038_s41598-024-51154-z]. Lightweight cryptographic protocols suitable for resource-constrained sensors, blockchain-based audit logs for immutable provenance trails, and anomaly detection monitoring for unusual data access patterns constitute a defense-in-depth security architecture necessary for protecting sensitive classroom data.
@@
Policy frameworks and standards alignment leverage guidelines and interoperability practices proven in Active Assisted Living (AAL) and health AmI contexts, including policy templates for consent, retention, data sharing agreements, and cross-institutional governance structures [DOI_2196_15923, DOI_3390_smartcities5030053, DOI_3390_en14072024]. Formal policy mechanisms for context-sensitive data acquisition and access gating developed for distributed cyber-physical systems can be repurposed for school district networks, specifying role-based access controls, purpose limitations, and automated policy enforcement [DOI_3390_computers10080101].
@@
Together, these techniques form a comprehensive privacy architecture layering hardware selection, local processing, algorithmic obfuscation, cryptographic protection, provenance tracking, and institutional governance to minimize identifiability while preserving pedagogically actionable signals [DOI_1109_tmc_2021_3092271, DOI_1007_s41019-020-00118-0, DOI_33166_aetic_2020_05_001, DOI_1109_access_2020_2995917, DOI_1038_s41598-024-51154-z].
@@ 3.14.6 Ethical Frameworks, Stakeholder Engagement, and Participatory Design
Early and iterative stakeholder engagement involves end users (students with developmentally appropriate assent processes), caregivers (parents/guardians providing informed consent), teachers (implementing technology and interpreting analytics), administrators (ensuring policy compliance and resource allocation), and independent ethics advisors throughout design cycles to surface values, acceptability thresholds, and perceived risks [DOI_1016_s2589-7500(20)30275-2, DOI_1016_j_jval_2021_11_1372, DOI_2196_47586]. MMLA with young children requires parent/guardian informed consent documents that clearly explain sensor modalities, data uses, retention policies, access controls, and opt-out procedures in accessible language, complemented by child-appropriate assent processes using visual aids and age-appropriate explanations respecting children’s developing autonomy [DOI_1111_bjet_12959, DOI_1016_s2589-7500(20)30275-2].
Values-oriented and participatory design methodologies, adapted from AmI healthcare platform development, align technology architectures with pedagogical aims, community norms, and ethical values through structured design activities [DOI_1016_j_jval_2021_11_1372, DOI_1155_2021_5518722, DOI_1016_s2589-7500(20)30275-2]. Co-design workshops with teachers and parents identify acceptable use cases, unacceptable surveillance practices, and preferred transparency mechanisms, while child-centered design activities (drawing desired classroom technologies, role-playing sensor scenarios) elicit children’s perspectives on privacy and technology in their learning environments.
Risk assessment and duty of care require formal privacy, psychosocial, and pedagogical risk assessments before deployment, recognizing that continuous behavioral monitoring may alter classroom relationships, create surveillance anxiety, produce stigmatizing inferences, or generate data vulnerable to misuse [DOI_1016_s2589-7500(20)30275-2, DOI_2196_47586, DOI_1001_jama_2019_21699]. AmI healthcare literature highlights how ambient sensing can subtly shift power dynamics and trust relationships, underscoring the need for ongoing risk monitoring, incident response protocols, and mechanisms for students/families to raise concerns without penalty.
Legal and institutional alignment ensures technical architectures comply with applicable education privacy laws (FERPA in U.S. contexts, GDPR in European contexts), district data governance policies, and ethical research standards when MMLA systems serve dual educational and research purposes [DOI_1001_jama_2019_21699, DOI_2196_15923]. Clear policies specifying data ownership (student/family, school district, technology vendor), retention limits (automated deletion schedules, prohibition on indefinite retention), permissible uses (restricted to specified pedagogical or research purposes), and prohibited uses (commercial profiling, predictive discipline, third-party sale) provide institutional safeguards complementing technical protections.
@@ 3.14.7 User Experience and Teacher Dashboard Design
UX evaluation and acceptability testing are essential components of AmI deployment, with standardized evaluation instruments available for measuring perceived usefulness, trust, privacy concerns, cognitive load, and unintended consequences [DOI_1155_2021_5518722, DOI_1016_j_jval_2021_11_1372]. For elementary MMLA interfaces, design priorities emphasize presenting aggregated, actionable indicators to teachers (class-level engagement trends, zone-based attention heatmaps, activity-type participation patterns) rather than granular individual student judgments, unless explicit consent, demonstrated pedagogical necessity, and strict governance safeguards justify finer-grained displays [DOI_1111_bjet_12959, DOI_1016_j_jval_2021_11_1372, DOI_1016_s2589-7500(20)30275-2].
Teacher dashboards should support pedagogical decision-making without overwhelming educators with excessive detail or creating punitive surveillance atmospheres. Effective designs provide at-a-glance class climate indicators (overall engagement level, attention distribution across learning zones, participation balance), customizable alerts for situations warranting intervention (prolonged disengagement cluster, environmental conditions degrading below thresholds), and historical trend visualizations supporting reflection on instructional effectiveness [DOI_1111_bjet_12959]. Critically, dashboards must avoid reducing complex learning processes to simplistic numeric scores or traffic-light color schemes that encourage superficial interpretations or stigmatizing labels.
@@ 3.14.8 Operational Blueprint: Stepwise Implementation Checklist
Phase 1 - Co-design and needs analysis: Convene teachers, parents, school administrators, and child representatives using structured participatory design methods; define pedagogical objectives, acceptable intervention modes, and privacy boundaries grounded in values-oriented design frameworks [DOI_1111_bjet_12959, DOI_1016_j_jval_2021_11_1372, DOI_1016_s2589-7500(20)30275-2].
@@
Phase 2 - Risk assessment and policy specification: Conduct formal privacy impact assessment, psychosocial risk analysis, and pedagogical necessity review; specify data retention schedules, consent procedures, access control matrices, incident response protocols, and audit mechanisms modeled on AAL and AmI healthcare guidance [DOI_2196_15923, DOI_1016_s2589-7500(20)30275-2, DOI_2196_47586].
@@
Phase 3 - Sensor selection and pilot hardware: Choose least-identifying sensor modalities meeting pedagogical requirements (e.g., low-resolution thermal arrays for occupancy, capacitive floors for trajectories, environmental sensors for context, aggregated device logs for participation) while avoiding routine RGB camera or continuous audio use unless explicitly justified and safeguarded [DOI_3390_s21041036, DOI_1109_jsen_2021_3139442, DOI_3233_ais-220577, DOI_3390_electronics10050559].
@@
Phase 4 - Edge-centric architecture and secure networking: Design classroom edge computing nodes for local feature extraction and obfuscation, implement secure key management and encrypted communication channels, deploy intrusion detection monitoring, and configure network policies to transmit only aggregated features as specified by governance policies [DOI_33166_aetic_2020_05_001, DOI_1109_access_2020_2995917, DOI_1038_s41598-024-51154-z].
@@
Phase 5 - Algorithmic design and privacy measures: Implement sensor-appropriate machine learning models (CNNs for spatial sensor fusion, LSTMs for temporal trajectory modeling) with feature-level obfuscation techniques, differential privacy mechanisms where applicable, and minimal necessary data retention enforced through automated deletion schedules [DOI_3390_s21041036, DOI_3390_electronics10121390, DOI_1109_tmc_2021_3092271, DOI_1155_2020_8876782].
@@
Phase 6 - Provenance, auditing, and interoperability: Tag all processed outputs with provenance metadata documenting origin, processing steps, consent status, and retention policies to enable auditing and regulatory compliance; adopt interoperable metadata schemas and ontology practices to support potential future integration while preserving policy-driven access controls [DOI_1007_s41019-020-00118-0, DOI_3390_smartcities5030053, DOI_3390_en14072024].
@@
Phase 7 - Pilot evaluation and UX testing: Deploy system in small-scale classroom pilots using standardized UX questionnaires and pedagogical outcome measures; evaluate teacher workflow impacts, psychosocial effects on students and classroom climate, technical reliability, and model performance; iterate design based on stakeholder feedback [DOI_1155_2021_5518722, DOI_1016_j_jval_2021_11_1372, DOI_1111_bjet_12959].
@@
Phase 8 - Scale-up governance and continuous monitoring: Establish ongoing risk monitoring processes, model maintenance and retraining schedules to address concept drift, stakeholder feedback channels for reporting concerns, periodic audits of data access logs and policy compliance, and governance committee oversight with teacher, parent, and administrator representation [DOI_3390_s23187729, DOI_1016_s2589-7500(20)30275-2, DOI_1016_j_ress_2022_108889].
@@ 3.14.9 Minimal-Risk Reference Configuration
Sensors: Overhead low-resolution thermal array (8x8 or 16x16 pixels) for non-identifying occupancy and zone activity classification [DOI_3390_s21041036]; capacitive proximity floor tiles for trajectory extraction without visual capture [DOI_1109_jsen_2021_3139442]; ambient environmental sensors (CO2, noise level, illuminance) for contextual class-level activity proxies [DOI_3233_ais-220577]; aggregated classroom device event logs from tablets/interactive displays for participation metrics [DOI_33166_aetic_2020_05_001].
Architecture: Classroom edge computing node executes CNN-based occupancy inference and LSTM-based short-horizon trajectory summarization, applies feature-level obfuscation via camouflage learning techniques, and exports only zone-aggregated engagement scores and participation summaries to teacher dashboard; raw sensor streams retained transiently on edge device (minutes to hours) for debugging purposes per policy, then automatically purged [DOI_3390_s21041036, DOI_3390_electronics10121390, DOI_1109_tmc_2021_3092271, DOI_33166_aetic_2020_05_001].
Governance: Informed parental/guardian consent with child assent; teacher dashboard limited to class/zone-level aggregates without individual student tracking; provenance labels attached to all exported data summaries documenting consent status and permissible uses; regular independent audits of data access logs; family opt-out pathway allowing non-participation without penalty [DOI_1016_s2589-7500(20)30275-2, DOI_1016_j_jval_2021_11_1372, DOI_1007_s41019-020-00118-0].
This configuration follows AmI precedents for non-identifying sensing and edge-centric processing to preserve privacy while enabling pedagogically valuable real-time engagement awareness [DOI_3390_s21041036, DOI_1109_jsen_2021_3139442, DOI_33166_aetic_2020_05_001, DOI_1109_tmc_2021_3092271].
@@ 3.14.10 Persistent Challenges and Research Gaps
Limited elementary evidence base: Systematic reviews of MMLA with young children identify comparatively small sample sizes, methodological heterogeneity, and limited longitudinal validation, indicating urgent need for domain-specific validation studies employing child-centered evaluation frameworks and developmentally appropriate measures [DOI_1111_bjet_12959].
Model interpretability and fairness: Adoption of computer vision and affective computing in educational contexts demands heightened caution regarding over-interpretation of algorithmic inferences, fairness across demographic subgroups, and risks of misclassification producing stigmatizing labels or inappropriate interventions; medical CV literature emphasizes the necessity of interpretability, domain-specific validation, and human oversight when algorithms influence consequential decisions about children [DOI_1038_s41746-020-00376-2].
System sustainability and model lifecycle: HAR and AmI models exhibit finite operational lifespans due to concept drift, environmental changes, and evolving classroom practices, requiring robust maintenance plans, continuous performance monitoring, and periodic retraining protocols to sustain accuracy and prevent silent degradation [DOI_3390_s23187729].
Interoperability and metadata standards: While ontologies and metadata schemas exist for smart buildings and cyber-physical systems, educational technology ecosystems lack widely adopted standards for MMLA data exchange, creating barriers to cross-institutional research, longitudinal student tracking across grade transitions, and integration with existing learning management systems [DOI_3390_smartcities5030053, DOI_3390_en14072024, DOI_3390_su13105578].
IoT security and attack surfaces: Classroom IoT sensors present unique security challenges including physical tampering risks, wireless network vulnerabilities, and resource constraints limiting cryptographic capabilities, demanding standardized secure provisioning frameworks and tailored intrusion detection systems [DOI_1038_s41598-024-51154-z, DOI_1109_access_2020_2995917, DOI_1109_access_2022_3179418].